I am getting 30 baby chickens. I am going to keep them in a brooder that consist of a cardboard roll that has been rolled out into a circle. The only thing is I'm not sure how big to make the circle. I read that there should be 1/2 a square foot per chick. What would the area and diameter be of this circle?
Found 2 solutions by stanbon, scott8148:
Answer by stanbon(75887)

I am getting 30 baby chickens. I am going to keep them in a brooder that consist of a cardboard roll that has been rolled out into a circle. The only thing is I'm not sure how big to make the circle. I read that there should be 1/2 a square foot per chick. What would the area and diameter be of this circle?
----
30 chicks would need 30(1/2) = 15 sq yrds of area
------
Area of a circle is (pi)r^2
---
Solve:
(pi)r^2 = 15
r^2 = 15/(pi)
r = sqrt(4.7746) = 2.1851 yrds.
-----
Diameter = 2(2.1851) = 4.36 yrds.
======================================
Cheers,
Stan H.
================

Answer by scott8148(6628)

30 chicks at 1/2 ft^2 each is 15 ft^2

pi * d^2 / 4 = 15

d^2 = 60 / pi ___ d = sqrt(60 / pi)